OpenStudy (anonymous):

PT tan (π/4 + θ) – tan (π/4 –θ) = 2 tan ^2 θ

OpenStudy (experimentx):

what's PT?

OpenStudy (anonymous):

\[\tan (A+B) =\frac{ tanA+tanB }{1-tanAtanB}\]

OpenStudy (anonymous):

PT stands for Prove That

OpenStudy (anonymous):

msingh , apply that formula for tan (π/4 + θ)

OpenStudy (anonymous):

\[\tan(A-B) = \frac{ tanA - tanB }{1+tanAtanB }\]

OpenStudy (anonymous):

apply this formula for tan (π/4 - θ)

OpenStudy (anonymous):

Simplify it~

OpenStudy (anonymous):

let me know if u have any doubts further?

OpenStudy (anonymous):

i m getting 4tan(theta) /(1-tan^theta)

OpenStudy (anonymous):

@SandeepReddy

OpenStudy (anonymous):

Yes msingh

OpenStudy (anonymous):

After solving i m getting 4tan(theta) /(1-tan^theta)

OpenStudy (anonymous):

can anybody tell after this

OpenStudy (anonymous):

k @experimentX thank u

OpenStudy (experimentx):

yw

OpenStudy (anonymous):

@experimentX what do u mean by "the identity is false"?

OpenStudy (experimentx):

the left hand side is not equal to the right hand side.

OpenStudy (anonymous):

Yeah @msingh must be writing it \[\tan (π/4 + θ) – \tan (π/4 –θ) = 2 \tan 2 θ\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

So u clear about it msingh?

OpenStudy (anonymous):

so how to solve next

OpenStudy (anonymous):

from 4tan(theta) /(1-tan^theta) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

from 4tan(theta) /(1-tan^2theta) ?

OpenStudy (anonymous):

\[\frac{ 4\tan{\theta} }{ 1-\tan^2{\theta} }=2*\frac{ 2\tan{\theta} }{ 1-\tan^2{\theta} }\] and we have a relation \[\tan{2\theta} = \tan({\theta+\theta}) =\frac{ \tan{\theta}+\tan{\theta} }{ 1-\tan{\theta}*\tan{\theta} }=\frac{ 2\tan{\theta} }{ 1-\tan^2{\theta} }\] therefore you have \[\frac{ 4\tan{\theta} }{ 1-\tan^2{\theta} } = 2\tan{2\theta}\]

OpenStudy (anonymous):

@SandeepReddy thank yu so much

OpenStudy (anonymous):

My pleasure! ( and bit pain while writing equations though! :p )